Iola has $75. She buys a pair of shoes on sale for one-half off and a pair of socks for $6. She has $32 left. Which equation can be used to find x, the regular price of the shoes?

Answer by ewatrrr(24785)

x the regular price of shoes
$75 - x/2 - $6 = $32
75 - 38 = x/2
37 = x/2
$74 = x, the regular price of shoes
And...checking
$75 - $37-$6 = $32